OpenStudy (anonymous):

Solve the equation by completing the square x^2-4x-22=0

OpenStudy (anonymous):

x^2-4x=22

OpenStudy (anonymous):

take the middle term -4 and divide it by 2 which is -4/2=-2 and then square this to get (-2)^2=4 add this to both sides of the equation yielding x^2-4x+4=22+4 the left hand side is now a perfect trinomial square and can be factored down to (x-2)^2=26

OpenStudy (anonymous):

sqrt both sides (x-2)= +/-sqrt(26) then add the 2 over to get x = 2+/-sqrt(26)

OpenStudy (anonymous):

a common mistake made during the use of this method is forgetting the +/- in front of the square root.

OpenStudy (anonymous):

i need a smaller value and a larger value of the equation

OpenStudy (anonymous):

your answers are 2+sqrt(26) and 2-sqrt(26)

OpenStudy (anonymous):

what is 26 squared?